# Upper bounds on the highest phonon frequency and superconducting temperature from fundamental physical constants

### v2

@@ 1 Introduction
Here, we show that fundamental physical constants provide an upper bound to the frequency of phonons in condensed matter phases, setting a limit to the vibrational frequency of atoms. This bound involves the dimensionless electron-to-proton mass ratio $\beta$, the key constant involved in setting the finely-tuned “habitable zone” of our world in the ($\alpha$, $\beta$) space, where $\alpha$ is the fine-structure constant [1]. We find this bound to be in agreement with ab initio simulations of atomic hydrogen and hydride superconductors which, as hydrogen is the lightest nucleus (a single proton), host the highest possible phonon frequencies. We subsequently show that this upper bound in the maximal phonon frequency implies an upper bound to the phonon-mediated superconducting critical temperature $T_{\mathrm{c}}$ in condensed matter. In particular, we show that fundamental constants set the upper bound to $T_{c}$ on the order of $10^{2}-10^{3}$ K, consistent with calculations of $T_{c}$ from optimal Eliashberg functions. We also observe that in addition to electron-phonon coupling, the bound to $T_{c}$ in terms of fundamental constants applies to other coupling mechanisms such as electron-electron coupling. This implies that the current search for $T_{\mathrm{c}}$ at and above $300$ K is itself due to fundamental constants taking their observed values. We finally observe that fundamental constants affect the observability and operation of other phenomena in condensed matter physics including phase transitions. This has wider implications compared to earlier work discussing how fundamental constants set bounds on properties that are already observed [15, 16, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13, 31, 12, 14].
@@ 2 Upper bound on the vibrational frequency
The maximal frequency in a given system, $\omega_{\mathrm{m}}$, is the largest frequency in the phonon dispersion, and is typically close to the Debye frequency $\omega_{\rm D}$. We are interested in an upper bound, $\omega_{\mathrm{m}}^{\mathrm{u}}$, on this maximal frequency (and thus an upper bound on all phonon frequencies). We recall two important properties of condensed matter phases: the interatomic separation $a$ and a cohesive, or bonding, energy $E$. The Bohr radius, $a_{\rm B}$, sets the characteristic scale of $a$ to the order of Angstroms:
@@
The ratio $\frac{\hbar\omega_{\mathrm{m}}}{E}$ can be derived by approximating $\hbar\omega_{\mathrm{m}}$ as $\hbar\left(\frac{E}{ma^{2}}\right)^{\frac{1}{2}}$, where $m$ is the atom mass and then either (a) using $a=a_{\rm B}$ from (1) and $E=E_{\rm R}$ from (2) or (b) using $E=\frac{\hbar^{2}}{2m_{e}a^{2}}$. This gives, up to a factor of order one, the following relation:
@@
Equations  (4)-(6) show that fundamental constants set $\omega_{\rm m}^{\rm u}$ on the order of $10^{3}$$10^{4}$ K, a result we will use to discuss phonon-mediated superconductivity below. The upper bound $\omega_{\rm m}^{\rm u}$ in (6) is of the same order of magnitude, but above, the experimental Debye frequency $\omega_{\rm D}$ of 2240 K in ambient pressure diamond [33], the hardest known material [34] or a similar value of the maximal phonon frequency in graphene [35].
@@
Before calculating $\omega_{\rm m}$, we first evaluate the effect of pressure on the fundamental bound (4) because solid atomic hydrogen exists at high pressure only as mentioned above. We make two remarks in relation to the pressure effect. First, hydrogen is a unique element with no core electrons. As compared to heavier elements, this gives weaker repulsive contributions to the interatomic interaction and weaker pressure dependence of elastic moduli [39]. As a result, pressure has a weaker effect on $\omega_{\rm m}$ in hydrogen as compared to other systems. Second, the effect of pressure on $\omega_{\rm m}^{\rm u}$ can be approximately estimated by adding the term $PV$ to $E$ in Eq. (3), representing the work needed to overcome the elastic deformation [40] due to external pressure in order to surmount the energy of cohesion. Here, $V$ is the elementary volume on the order of $a^{3}$. According to Eqs. (3)-(4), this increases $\omega_{m}^{u}(P)$ as
$\omega_{\rm m}^{\rm u}(P)=\omega_{\rm m}^{\rm u}+a^{3}\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}P$ (7)
@@
Figure 2: Maximal phonon frequency as a function of pressure for different systems. Blue bullet points show the calculated maximal phonon frequencies in atomic hydrogen at the corresponding pressures. Red diamonds show the maximal phonon frequencies in a range of hydrides at different pressures, which are, from left to right: MgIrH6 [48], MgVH6 [49], CaH6 [50], LaH10 (top) [51], H3S (bottom) [52], and SrH10 [53]. The horizontal solid black line shows the fundamental bound $\omega_{\rm m}^{\rm u}$ (6). The horizontal dash-dotted line shows the Debye frequency of diamond at ambient pressure. The dashed line on top of the graph shows $\omega_{\rm m}^{\rm u}(P)$ in Eq. (7) with $a=2a_{\rm B}$.
@@ 3.1 Upper limit to $T_{c}$
$T_{c}^{u}$ in Eq. (8) can be made more stringent in a particular family of materials. Recall that (a) $E_{\rm R}$ used to evaluate the bonding energy $E$ is larger than bonding energies in real materials and (b) the average mass in real materials is larger than the proton mass $m_{p}$ featuring in Eqs. (4) and (8). Both effects reduce $T_{c}^{u}$. According to Eqs. (3), (4), and (8), the difference between $E$ and $E_{\rm R}$ results in $T_{c}^{u}$ decreasing by a factor $\frac{E_{\rm{eff}}}{E_{\rm R}}$, where $E_{\rm{eff}}$ is the effective bonding energy in the relevant subsystem. The difference between the average mass and $m_{p}$ reduces $T_{c}^{u}$ by a factor $\left(\frac{m_{\rm{eff}}}{m_{p}}\right)^{\frac{1}{2}}$, where $m_{\rm{eff}}$ is the effective mass in the relevant subsystem. The same reasoning applies to the upper frequency bound (4).
@@
We also note that our bound (8) invited a recent discussion in Ref. [74]. This work derived an upper bound to $T_{c}$ using different arguments and interestingly showed that this bound is expressed as the same combination of fundamental constants as Eq. (8).
@@ Acknowledgements.
Appendix
.1 Density functional theory calculations
In our density functional theory calculations, we use the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [78], an energy cutoff of $1200$ eV and a $\mathbf{k}$-point grid of spacing $2\pi\times 0.025$-1 to sample the electronic Brillouin zone. We relax cell parameters and internal coordinates to obtain a pressure to within $10^{-4}$ GPa of the target pressure and forces smaller than $10^{-5}$ eV/Å. We calculate the phonon dispersion using the finite displacement method [79] in conjunction with nondiagonal supercells [80] with a $4\times 4\times 4$ coarse $\mathbf{q}$-point grid to sample the vibrational Brillouin zone, and we use Fourier interpolation to calculate the phonon frequencies along a high symmetry path in the Brillouin zone.
.2 Optimization of the Eliashberg function

### v3

@@ 1 Introduction
Here, we show that fundamental physical constants provide an upper bound to the frequency of phonons in condensed matter phases, setting a limit to the vibrational frequency of atoms in these phases. This bound involves the dimensionless electron-to-proton mass ratio $\beta$, the key constant involved in setting the finely-tuned “habitable zone” of our world in the ($\alpha$, $\beta$) space, where $\alpha$ is the fine-structure constant [1]. The bound applies to non-molecular systems, and we find it to be in agreement with ab initio simulations of atomic hydrogen and hydride superconductors which, as hydrogen is the lightest nucleus (a single proton), host the highest possible phonon frequencies. We subsequently show that this upper bound in the maximal phonon frequency implies an upper bound to the phonon-mediated superconducting critical temperature $T_{\mathrm{c}}$ in condensed matter. In particular, we show that fundamental constants set the upper bound to $T_{c}$ on the order of $10^{2}-10^{3}$ K, consistent with calculations of $T_{c}$ from optimal Eliashberg functions. We also observe that in addition to electron-phonon coupling, the bound to $T_{c}$ in terms of fundamental constants applies to other coupling mechanisms such as electron-electron coupling. This implies that the current search for $T_{\mathrm{c}}$ at and above $300$ K is itself due to fundamental constants taking their observed values. We finally observe that fundamental constants affect the observability and operation of other phenomena in condensed matter physics including phase transitions. This has wider implications compared to earlier work discussing how fundamental constants set bounds on properties that are already observed [15, 16, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 13, 31, 12, 14].
2 Methods
In our density functional theory calculations, we use similar methods to our earlier paper [28]. This includes using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [32], an energy cutoff of $1200$ eV and a $\mathbf{k}$-point grid of spacing $2\pi\times 0.025$-1 to sample the electronic Brillouin zone.
We relax cell parameters and internal coordinates to obtain a pressure to within $10^{-4}$ GPa of the target pressure and forces smaller than $10^{-5}$ eV/Å. We calculate the phonon dispersion using the finite displacement method [33] in conjunction with nondiagonal supercells [34] with a $4\times 4\times 4$ coarse $\mathbf{q}$-point grid to sample the vibrational Brillouin zone. We use Fourier interpolation to calculate the phonon frequencies along a high symmetry path in the Brillouin zone.
@@ 3 Upper bound on the vibrational frequency
The maximal frequency in a given condensed system, $\omega_{\mathrm{m}}$, is the largest frequency in the phonon dispersion, and is typically close to the Debye frequency $\omega_{\rm D}$. We are interested in an upper bound, $\omega_{\mathrm{m}}^{\mathrm{u}}$, on this maximal frequency (and thus an upper bound on all phonon frequencies). We recall two important properties of condensed matter phases: the interatomic separation $a$ and a cohesive, or bonding, energy $E$. The Bohr radius, $a_{\rm B}$, sets the characteristic scale of $a$ to the order of Angstroms:
@@
The ratio $\frac{\hbar\omega_{\mathrm{m}}}{E}$ can be derived by approximating $\hbar\omega_{\mathrm{m}}$ as $\hbar\left(\frac{E}{ma^{2}}\right)^{\frac{1}{2}}$, where $m$ is the atom mass and then either (a) using $a=a_{\rm B}$ from (1) and $E=E_{\rm R}$ from (2) or (b) using $E=\frac{\hbar^{2}}{2m_{e}a^{2}}$. This gives, up to a factor of order one ($\sqrt{2}$), the following relation:
@@
The upper bound $\omega_{\rm m}^{\rm u}$ in (6) is of the same order of magnitude, but above, the experimental Debye frequency $\omega_{\rm D}$ of 2240 K in ambient pressure diamond [36], the hardest known material [37] or a similar value of the maximal phonon frequency in graphene [38].
@@
Before calculating $\omega_{\rm m}$, we first evaluate the effect of pressure on the fundamental bound (4) because solid atomic hydrogen exists at high pressure only as mentioned above. We make two remarks in relation to the pressure effect. First, hydrogen is a unique element with no core electrons. As compared to heavier elements, this gives weaker repulsive contributions to the interatomic interaction and weaker pressure dependence of elastic moduli [42]. As a result, pressure has a weaker effect on $\omega_{\rm m}$ in hydrogen as compared to other systems. Second, the effect of pressure on $\omega_{\rm m}^{\rm u}$ can be approximately estimated by adding the term $PV$ to $E$ in Eq. (3), representing the work needed to overcome the elastic deformation [43] due to external pressure in order to surmount the energy of cohesion. Here, $V$ is the elementary volume on the order of $a^{3}$. According to Eq. (3), this increases $\omega_{m}^{u}(P)$ as
$\omega_{\rm m}^{\rm u}(P)=\omega_{\rm m}^{\rm u}+\left(\frac{m_{e}}{m_{p}}\right)^{\frac{1}{2}}\frac{a^{3}P}{\hbar}$ (7)
@@
Figure 2: Maximal phonon frequency as a function of pressure for different systems. Blue bullet points show the calculated maximal phonon frequencies in atomic hydrogen at the corresponding pressures. Red diamonds show the maximal phonon frequencies in a range of hydrides at different pressures, which are, from left to right: MgIrH6 (1) [51], MgVH6 (2) [52], CaH6 (3) [53], H3S (4) [54], LaH10 (5) [55] and SrH10 (6) [56]. The horizontal solid black line shows the fundamental bound $\omega_{\rm m}^{\rm u}$ (6). The horizontal dash-dotted line shows the Debye frequency of diamond at ambient pressure. The dashed line on top of the graph shows $\omega_{\rm m}^{\rm u}(P)$ in Eq. (7) with $a=2a_{\rm B}$.
@@
Our derivation of the upper frequency bound is mostly suited to evaluate the highest oscillation frequency of acoustic phonon modes or in systems with no substantial hierarchy of bonding energies and frequencies (recall that we apply Eq. (2) for the bonding energies in the system). On the other hand, the intra-molecular frequencies in molecular solids (“vibrons”) can be considerably higher than the rest of the frequency spectrum, and the highest oscillation frequency can be higher than $\omega_{\rm m}^{\rm u}$ in these systems. Considering, for example, H2 molecule and using the reduced mass $m/2$ increases $\omega_{\rm m}$ in Eq. (3) and $\omega_{\rm m}^{\rm u}$ in Eq. (6) by $\sqrt{2}$. Recalling that the approximate derivation of Eq. (3) dropped $\sqrt{2}$ on the right-hand side, the highest oscillation frequency in hydrogen-containing molecules and molecular solids can increase by a factor of 2. Accordingly, one could choose to define the upper bound for the oscillation frequency as $73602\omega_{\rm m}^{\rm u}=$ 7360 K, in which case it would include molecules and molecular solids: highest intra-molecular frequencies are about 6200-6400 K and lower than $2\omega_{\rm m}^{\rm u}$. This could be of interest from the point of view of having a maximal oscillation frequency in all systems including molecular ones. However, we note that in view of the approximations involved in Eqs. (3) and (4), one of the main purposes of Eq. (6) is to evaluate the order of magnitude of the maximal frequency from fundamental constants only.
In addition to deriving the upper frequency bound, the two main results in this section are: (a) the upper bound applies well to non-molecular systems and (b) Eqs.  (4) and (6) show that fundamental constants set $\omega_{\rm m}^{\rm u}$ on the order of $10^{3}$$10^{4}$ K, the result we will use to discuss phonon-mediated superconductivity below.
@@ 4.1 Upper limit to $T_{c}$
$T_{c}^{u}$ in Eq. (8) can be made more stringent and adaptable to a particular family of materials. Recall that (a) $E_{\rm R}$ used to evaluate the bonding energy $E$ is larger than bonding energies in real materials and (b) the average mass in real materials is larger than the proton mass $m_{p}$ featuring in Eqs. (4) and (8). Both effects reduce $T_{c}^{u}$. According to Eqs. (3), (4), and (8), the difference between $E$ and $E_{\rm R}$ results in $T_{c}^{u}$ decreasing by a factor $\frac{E_{\rm{eff}}}{E_{\rm R}}$, where $E_{\rm{eff}}$ is the effective bonding energy in the relevant subsystem. The difference between the average mass and $m_{p}$ reduces $T_{c}^{u}$ by a factor $\left(\frac{m_{\rm{eff}}}{m_{p}}\right)^{\frac{1}{2}}$, where $m_{\rm{eff}}$ is the effective mass in the relevant subsystem. The same approximate reasoning applies to the upper frequency bound (4).
@@
We also note that our bound (8) invited a recent discussion in Ref. [77]. This work derived an upper bound to $T_{c}$ using a different type of arguments and interestingly showed that this bound is expressed as the same combination of fundamental constants as Eq. (8).
@@ Acknowledgements.
Appendix. Optimization of the Eliashberg function